OpenStudy (anonymous):

Rationalize the denominator and simplify. a) 5/4- √3 b) 12/√7 + 2 c) 6/√2 - √3 d) 15√6/3√5

OpenStudy (anonymous):

To rationalize denominators containing radicals, typically you multiply by the conjugate over itself (essentially multiplying by one). Example: a) \[\frac{ 5 }{ 4-\sqrt{3} } \times \frac{ 4+\sqrt{3} }{ 4+\sqrt{3} }\]